OpenStudy (anonymous):

Find the product. (-4·3·2)^2

OpenStudy (sandra):

go ahead and do the inside first if you want to: -4*3*2 = 24. -24^2 =?

OpenStudy (sandra):

(-7p)^3 = -7^3p^3 = (-7*-7*-7) * p^3 = ?

OpenStudy (anonymous):

=(-24)^2=576